Disjoint subsets A and B it a universal set U, except assume that n(U)=75, n(A)=40, and n(B)= 15. Then n(A'(complement)union B)=?
Answer by ikleyn(52781)

.
Why do not write the condition accurately and correctly, in order for the post itself and each word inside would make sense ????????



------------
Comment from student: Disjoint subsets A and B are in a universal set U, assume that
the number  n(U)=75,  the number n(A)=40, and the number n(B)= 15.  Then what is the number for  (A' ∪ B) ?
------------



My response:  In this case,  the subset  B is contained in  A' and is the part of  A'.

                      Therefore,  n(A' U B) = n(A') = n(U) - n(A) = 75 - 40 = 35.


Solved and answered.


---------
It was a pleasure to work with the text that you sent to me this time.


I am very glad that you followed my advise.


Happy learning !!